OpenStudy (anonymous):

Chapter scale:How to find actual length or distance by using scale and also how to find length or distance on a map @mathmale @Auxuris

OpenStudy (shinalcantara):

Say the scale for the map is 1 calibration(unit) is equal to 1 mi.. then in finding the actual distance, you just have to convert the number of units to the equivalent value say:  then if you're asked for the actual distance for 4 units then you'll have \[4 units(\frac{ 1 mi }{ 1 unit }) = 4 mi\]

OpenStudy (anonymous):

ok thnks

OpenStudy (mathmale):

If you have an actual map in front of you, look for the "scale" of the map. This might read "1 inch represents 25 miles" or something like that.

OpenStudy (mathmale):

Just supposing that you measure the distance from one point on this map to another and find that it is 2.1 inches. What is the distance (in miles)?

OpenStudy (mathmale):

Method: \[\frac{ 2.1~inches }{ 1 }*\frac{ 25 miles }{ 1~inch }=?\]

OpenStudy (mathmale):

Note that you can cancel 'inches' from numerator and denominator.

OpenStudy (anonymous):

let's take it m and km please then i'm gonna more understand it

OpenStudy (mathmale):

If you have an actual map in front of you, look for the "scale" of the map. This might read "1 cm represents 10 Km" or something like that.

OpenStudy (mathmale):

Suppose you use a metric ruler to measure the distance (in cm) from one point on your map to another point. Suppose that that distance is 5.7 cm. What is the equivalent distance in Km?

OpenStudy (mathmale):

\[\frac{ 5.7 cm }{ 1 }*\frac{ 10~Km }{1~cm }\] cancel the unit "cm" and multiply 5.7 by 10.